# An Effective Flow-based Method for Positive-Unlabeled Learning: 2-HNC

### v1

@@ 3.2 Graph Notations
For an undirected graph $G=(V,E)$, where $V$ is the set of nodes (or vertices) and each edge $[i,j]\in E$ has a weight of $w_{ij}$, we define the function $C(V_{1},V_{2})$ for two sets $V_{1},V_{2}\subseteq V$: $C(V_{1},V_{2}):=\sum_{[i,j]\in E,i\in V_{1},j\in V_{2}}w_{ij}$. That is, this is the sum of weights of edges that have one endpoint in $V_{1}$ and the other in $V_{2}$.
@@
The minimum $(s,t)$-cut problem is defined on a directed graph $G=(V,A)$ in which two nodes in $V$ are specified to be the source node, $s$, and the sink node, $t$. We call the directed graph with two designated nodes $s$ and $t$ an $(s,t)$-graph. An $(s,t)$-cut is a partition of the set of nodes $V$ into $(S,T)$ such that $s\in S,t\in T,S\subseteq V$ and $T=V\backslash S$. The set $S$ in the partition is referred to as the source set whereas the set $T$ is the sink set. The capacity of the $(s,t)$-cut $(S,T)$ is $C(S,T)$. A minimum $(s,t)$-cut is the $(s,t)$-cut that has the minimum capacity among all cuts. For the remainder of this paper, we will refer to the minimum $(s,t)$-cut simply as the minimum cut.
@@ 3.4 Hochbaum’s Normalized Cut (HNC)
We represent a dataset as an undirected graph $G=(V,E)$, where each node in the set $V$ corresponds to a sample in the data. The notation $V$ refers henceforth to both the dataset and the set of nodes in the graph. An edge $[i,j]\in E$ connects a pair of nodes (or samples), carrying the weight $w_{ij}$ that quantifies the pairwise similarity between samples $i$ and $j$.
For a partition of the set of samples $V$ into two disjoint sets $S$ and $\overline{S}$, where $\overline{S}=V\backslash S$, $C(S,\overline{S})=\sum_{[i,j]\in E,i\in S,j\in\overline{S}}w_{ij}$ represents the inter-similarity between samples in $S$ and its complement $\overline{S}$. $C(S,S)=\sum_{[i,j]\in E,i\in S,j\in S,i<j}w_{ij}$ denotes the total sum of similarities between samples within the set $S$, the intra-similarity of $S$.
@@
One might consider a variant of HNC that incorporates both intra-similarities, $C(S,S)$ and $C(\overline{S},\overline{S})$, and as such generalizing both HNC+ (1) and HNC- (2). This variant, with two tradeoff weights $\alpha\geq 0$ and $\beta\geq 0$, is given as problem (HNC±$\pm$):
As proved in the next theorem, HNC±$\pm$ (3) is equivalent to either HNC+ or HNC-, depending on the relative values of $\alpha$ and $\beta$.
@@
HNC±$\pm$ (3) is equivalent to HNC+ (1) when $\alpha\geq\beta$ for $\mu=\frac{\alpha-\beta}{1+\beta}$, and is equivalent to HNC- (2) when $\alpha<\beta$ for $\mu=\frac{\beta-\alpha}{1+\alpha}$.
@@ 4.1.2 Stage 2: Solving HNC+ with Positive Labeled Samples and Likely-Negative Unlabeled Samples
SelectNeg takes as input the sequence of optimal data partitions $(S^{*}_{1},\overline{S}^{*}_{1}),(S^{*}_{2},\overline{S}^{*}_{2}),\dots,(S^{*}_{q},\overline{S}^{*}_{q})$, which are the results of solving HNC- for a list of nonnegative and increasing values of $\mu$. The nested sequence $\overline{S}^{*}_{1}\subseteq\overline{S}^{*}_{2}\subseteq\dots\subseteq\overline{S}^{*}_{q}$ starts from $\overline{S}^{*}_{1}=\varnothing$ and expands until $\overline{S}^{*}_{q}=V\backslash L^{+}$, which is the largest possible since we require $L^{+}$ to be in $S_{q}^{*}$.
@@
Once the likely-negative set $L^{N}$ is formed, we use the positive labeled set $L^{+}$ and the likely-negative set $L^{N}$ as the seed sets for $S$ and $\overline{S}$, respectively, and solve HNC+ with the seed set constraint $L^{+}\subseteq S\subseteq V\backslash L^{N}$. As a result, the output from stage 2 is another sequence of data partitions, which are the optimal solutions to HNC+ for the list of values of the tradeoff parameter $\mu$.
@@ 4.2.1 Solving HNC+ and HNC- for a Tradeoff Parameter $\mu$ as a Minimum Cut Problem
In a binary classification problem where both labeled sets $L^{+}$ and $L^{-}$ are given, the seed set constraint is $L^{+}\subseteq S\subseteq V\backslash L^{-}$. The remaining samples in $V\backslash(L^{+}\cup L^{-})$ make up the unlabeled set $U$. Under this constraint, the solution to HNC+ for a tradeoff parameter $\mu$, which is solved as (4) with a tradeoff parameter $\lambda=\frac{\mu}{\mu+2}$, is obtained from the minimum cut solution of the associated $(s,t)$-graph, $G^{+}_{st}(\lambda)$. The use of the superscript “$+$” for $G^{+}_{st}(\lambda)$ is to reflect the emphasis on the intra-similarity of $S$, or the positive prediction set, in HNC+.
The graph $G^{+}_{st}(\lambda)$ associated with (4) with the constraint $L^{+}\subseteq S\subseteq V\backslash L^{-}$ is illustrated in Figure 2(a). The graph is constructed as follows. The nodes in the graph consist of the set of nodes that represent samples in $V$, and two appended nodes: the source node $s$ and the sink node $t$. For each pair of samples $i,j\in V$, there are arcs $(i,j)$ and $(j,i)$, whose capacities are equal to the pairwise similarity between $i$ and $j$, $w_{ij}$. The set of source adjacent arcs is $A_{s}:=\{(s,i)|i\in L^{+}\cup U\}$ where the capacities of $(s,i)$ for $i\in L^{+}$ is $\infty$, and for $i\in U$, is $\lambda d_{i}$. The set of sink adjacent arcs is $A_{t}:=\{(i,t)|i\in L^{-}\}$. Each $(i,t)$ has a capacity of $\infty$.
It was shown by Hochbaum (2010) that, if ($(\{s\}\cup S^{*},\{t\}\cup\overline{S}^{*})$ is the minimum cut solution of $G^{+}_{st}(\lambda)$, then ($S^{*},\overline{S}^{*})$ is the optimal solution to HNC+. The proof provided in (Hochbaum, 2010) is omitted here. Accordingly, we predict that unlabeled samples in $S^{*}$ are positive, and those in $\overline{S}^{*}$ negative.
@@
the constraint $L^{+}\subseteq S\subseteq V\backslash L^{-}$.
@@
the constraint $L^{+}\subseteq S\subseteq V\backslash L^{-}$.
@@
In PU learning, negative labeled samples are not given. Therefore, $L^{-}=\varnothing$. HNC+ and HNC- in this context are then solved, for a tradeoff parameter $\lambda$, as minimum cut problems on the graphs in Figure 3(a) and 3(b), which are $G_{st}^{+}(\lambda)$ and $G_{st}^{-}(\lambda)$ where $L^{-}=\varnothing$. As explained in Section 4.1.1, HNC+, with $L^{-}=\varnothing$, has a trivial solution for all $\lambda\geq 0$. This is also reflected in the minimum cut of $G_{st}^{+}(\lambda)$ (Figure 3(a)) with $L^{-}=\varnothing$, which is ($(\{s\}\cup V,\{t\})$, as $t$ is disconnected from other nodes. Hence, in stage 1, we solve only HNC- using the graph $G_{st}^{-}(\lambda)$ in Figure 3(b). Once the likely-negative samples are used as seed samples in stage 2 (Section 4.1.2), HNC+ can be solved using the graph $G_{st}^{+}(\lambda)$ in Figure 2(a).
@@ 4.2.2 Solving HNC+ and HNC- for All Tradeoff Values in a Given List with a Parametric Cut Procedure
Graphs $G_{st}^{+}(\lambda)$ and $G_{st}^{-}(\lambda)$ are parametric flow networks in that the capacities of source-adjacent and sink-adjacent arcs ($(s,i)$ and $(i,t)$ for $i\in V$) are monotone non-increasing and non-decreasing with the parameter $\lambda$, or vice versa. For instance, $G_{st}^{+}(\lambda)$ in Figure 2(a), has source-adjacent capacities that can only increase with $\lambda$, and sink-adjacent capacities that are fixed. Note that both graphs are also parametric flow graphs with the parameter $\mu$ since $\lambda=1-\frac{2}{\mu+2}$ is an increasing function with respect to $\mu$. The minimum cuts in a parametric flow network are solved for all values of the parameter in the complexity of a single minimum cut procedure using the parametric cut (flow) algorithm, (Gallo et al., 1989; Hochbaum, 1998, 2008). The first is based on the push-relabel algorithm, and the latter two on the HPF (pseudoflow) algorithm.
@@ 4.2.3 Sparsification
A general drawback of using minimum cut in dense graphs is that the solution tends to not favor “balanced" partitions. In a balanced partition, there is a constant fraction $f<1$ of nodes on one side, and the number of edges between the two sets in the partition is $fn(1-f)n$, which is $O(n^{2})$. In that case, even if many edges in the partition have small capacities, their sheer number makes the capacity of such cuts much higher than cuts that contain a small number of nodes on one side. In the graphs we study, all pairwise similarities are evaluated. Therefore, such graphs are complete and dense. One approach for addressing this issue is to apply graph sparsification. The sparsification technique that we found to be effective for 2-HNC is the $k$-nearest neighbor (kNN) sparsification (Blum and Chawla, 2001) where samples $i$ and $j$ are connected only if $i$ is among the $k$ nearest neighbors of $j$, or vice versa. This results in a graph representation $G=(V,E)$ where $E$ is the set of pairs of samples where in each pair, one sample is one of the $k$ nearest neighbors of the other sample in the pair.
@@ 5 Implementation of 2-HNC
This section includes the specification of several implementation details. First, we give a brief description of the parametric cut solver used in this work. Second, we describe the choice of $k$ in the k-nearest neighbor graph sparsification method, as mentioned in the previous section. Finally, we quantify the pairwise similarity measure between pairs of samples.
@@ 5.1 Parametric Cut Solver
The values of $\lambda$ that we use for both $G_{st}^{+}(\lambda)$ and $G_{st}^{-}$ are evenly spaced increments of 0.001, starting from 0 and ending at 0.500, yielding a total of 501 values: ${0,0.001,0.002,0.003,\dots,0.500}$. We select this range of $\lambda$ since it allows the sink set $\overline{S}$ to expand to $V\backslash L^{+}$, which is the largest set possible, in stage 1 and the source set to grow to $V\backslash L^{N}$, also largest possible, in stage 2 for all datasets in our experiments. The choice of the increment of $0.001$ is used since smaller increments do not produce additional partitions that are different from those generated by using the increment of $0.001$; hence, it is small enough for datasets used in the experiments. For other datasets that are not included here, the appropriate range and increment of values of $\lambda$ may vary depending on characteristics of the datasets such as the size and scale of feature values.
@@ 6.4 Results on Synthetic Datasets
The histogram plots in Figures 4(a) and 4(c) show that, on average, 2-HNC gives positive improvement in terms of classification accuracy over uPU and PU ET, as the red dashed lines (the mean) lie to the right of the black solid line (the zero line). The average improvements, reported in Table 3, show the improvements of 2-HNC over uPU and PUET by $5.33\%$ and $8.02\%$, respectively. The fractions of datasets where 2-HNC improves on these two benchmarks are more than half, and in particular, as compared to PU ET, 2-HNC produces better results on almost all datasets ($92.96\%$). However, 2-HNC does not deliver consistent improvement in accuracy over nnPU, with the average improvement that is slightly below zero, at $-0.44\%$, reported in Table 3 and shown in Figure 4(c) where the red dashed line (the mean) lies almost on the solid line (the zero line). The percentage of datasets where 2-HNC improves over nnPU is just above $50\%$.
@@ 6.5 Results on Real Datasets
This is also the case when we consider the balanced accuracy of 2-HNC and the best benchmark on MNIST where the p-value is $0.0711$, which is larger than $0.05$. Hence, it cannot be concluded with statistical significance that 2-HNC has a worse performance than nnPU. The underperformance of 2-HNC is statistically significant for its accuracy on Phishing and MNIST as compared to PU ET, in which the p-values are $0.0004$ and $-93.3e\text{-}$ 9, respectively.

### v2

@@ 3.2 Graph Notations
For an undirected graph $G=(V,E)$, where $V$ is the set of nodes (or vertices) and each edge $i,j]\in E$ has a weight of $w_{ij}$, we define the function $C(V_{1},V_{2})$ for two sets $V_{1},V_{2}\subseteq V$: $C(V_{1},V_{2}):=\sum_{[i,j]\in E,i\in V_{1},j\in V_{2}}w_{ij}$. That is, this is the sum of weights of edges that have one endpoint in $V_{1}$ and the other in $V_{2}$.
@@
[…]
@@ 3.4 Hochbaum’s Normalized Cut (HNC)
We represent a dataset as an undirected graph $G=(V,E)$, where each node in the set $V$ corresponds to a sample in the data. The notation $V$ refers henceforth to both the dataset and the set of nodes in the graph. An edge $i,j]\in E$ connects a pair of nodes (or samples), carrying the weight $w_{ij}$ that quantifies the pairwise similarity between samples $i$ and $j$.
For a partition of the set of samples $V$ into two disjoint sets $S$ and $\overline{S}$, where $\S\overline{S}=V\backslash$ S, $C(S,\overline{S})=\sum_{[i,j]\in E,i\in S,j\in\overline{S}}w_{ij}$ represents the inter-similarity between samples in $S$ and its complement $\overline{S}$. $C(S,S)=\sum_{[i,j]\in E,i\in S,j\in S,i<j}w_{ij}$ denotes the total sum of similarities between samples within the set $S$, the intra-similarity of $S$.
@@
$(\text{HNC+})\quad\operatorname*{minimize}_{\varnothing\subset S\subset V}\quad C(S,\overline{S})-\mu\;C(S,S)$ (1)
$(\text{HNC-})\quad\operatorname*{minimize}_{\varnothing\subset S\subset V}\quad C(S,\overline{S})-\mu\;C(\overline{S},\overline{S})$ (2)
@@
[…]
@@ 4.1.2 Stage 2: Solving HNC+ with Positive Labeled Samples and Likely-Negative Unlabeled Samples
[…]
@@
[…]
@@ 4.2.1 Solving HNC+ and HNC- for a Tradeoff Parameter $\mu$ as a Minimum Cut Problem
In a binary classification problem where both labeled sets $L^{+}$ and $L^{-}$ are given, the seed set constraint is $\LL^{+}\subseteq S\subseteq V\backslash L^{-}$. The remaining samples in $V\backslash(L^{+}\cup L^{-})$ make up the unlabeled set $U$. Under this constraint, the solution to HNC+ for a tradeoff parameter $\mu$, which is solved as (4) with a tradeoff parameter $\lambda=\frac{\mu}{\mu+2}$, is obtained from the minimum cut solution of the associated $(s,t)$-graph, $G^{+}_{st}(\lambda)$. The use of the superscript “++” for $G^{+}_{st}(\lambda)$ is to reflect the emphasis on the intra-similarity of $S$, or the positive prediction set, in HNC+.
The graph $G^{+}_{st}(\lambda)$ associated with (4) with the constraint $\LL^{+}\subseteq S\subseteq V\backslash L^{-}$ is illustrated in Figure 2(a). The graph is constructed as follows. The nodes in the graph consist of the set of nodes that represent samples in $V$, and two appended nodes: the source node $s$ and the sink node $t$. For each pair of samples $i,j\in V$, there are arcs $(i,j)$ and $(j,i)$, whose capacities are equal to the pairwise similarity between $i$ and $j$, $w_{ij}$. The set of source adjacent arcs is $A_{s}:=\{(s,i)|i\in L^{+}\cup U\}$ where the capacities of $(s,i)$ for $i\in L^{+}$ is $\infty$, and for $i\in U$, is $\lambda d_{i}$. The set of sink adjacent arcs is $A_{t}:=\{(i,t)|i\in L^{-}\}$. Each $(i,t)$ has a capacity of $\infty$.
It was shown by Hochbaum (2010) that, if $(\{s\}\cup S^{*},\{t\}\cup\overline{S}^{*})$ is the minimum cut solution of $G^{+}_{st}(\lambda)$, then ($S^{*},\overline{S}^{*})$ is the optimal solution to HNC+. The proof provided in (Hochbaum, 2010) is omitted here. Accordingly, we predict that unlabeled samples in $S^{*}$ are positive, and those in $\overline{S}^{*}$ negative.
@@
the constraint $\LL^{+}\subseteq S\subseteq V\backslash L^{-}$.
@@
the constraint $\LL^{+}\subseteq S\subseteq V\backslash L^{-}$.
@@
In PU learning, negative labeled samples are not given. Therefore, $L^{-}=\varnothing$. HNC+ and HNC- in this context are then solved, for a tradeoff parameter $\lambda$, as minimum cut problems on the graphs in Figure 3(a) and 3(b), which are $G_{st}^{+}(\lambda)$ and $G_{st}^{-}(\lambda)$ where $L^{-}=\varnothing$. As explained in Section 4.1.1, HNC+, with $L^{-}=\varnothing$, has a trivial solution for all $\lambda\geq 0$. This is also reflected in the minimum cut of $G_{st}^{+}(\lambda)$ (Figure 3(a)) with $L^{-}=\varnothing$, which is $(\{s\}\cup V,\{t\})$, as $t$ is disconnected from other nodes. Hence, in stage 1, we solve only HNC- using the graph $G_{st}^{-}(\lambda)$ in Figure 3(b). Once the likely-negative samples are used as seed samples in stage 2 (Section 4.1.2), HNC+ can be solved using the graph $G_{st}^{+}(\lambda)$ in Figure 2(a).
@@ 4.2.2 Solving HNC+ and HNC- for All Tradeoff Values in a Given List with a Parametric Cut Procedure
Graphs $G_{st}^{+}(\lambda)$ and $G_{st}^{-}(\lambda)$ are parametric flow networks in that the capacities of source-adjacent and sink-adjacent arcs $($s,i)(s,i) and $(i,t)$ for $i\in V$) are monotone non-increasing and non-decreasing with the parameter $\lambda$, or vice versa. For instance, $G_{st}^{+}(\lambda)$ in Figure 2(a), has source-adjacent capacities that can only increase with $\lambda$, and sink-adjacent capacities that are fixed. Note that both graphs are also parametric flow graphs with the parameter $\mu$ since $\lambda=1-\frac{2}{\mu+2}$ is an increasing function with respect to $\mu$. The minimum cuts in a parametric flow network are solved for all values of the parameter in the complexity of a single minimum cut procedure using the parametric cut (flow) algorithm, (Gallo et al., 1989; Hochbaum, 1998, 2008). The first is based on the push-relabel algorithm, and the latter two on the HPF (pseudoflow) algorithm.
@@ 4.2.3 Sparsification
A general drawback of using minimum cut in dense graphs is that the solution tends to not favor “balanced" partitions. In a balanced partition, there is a constant fraction $f<1$ of nodes on one side, and the number of edges between the two sets in the partition is fn(1f)nfn(1-f)n, which is $O(n^{2})$. In that case, even if many edges in the partition have small capacities, their sheer number makes the capacity of such cuts much higher than cuts that contain a small number of nodes on one side. In the graphs we study, all pairwise similarities are evaluated. Therefore, such graphs are complete and dense. One approach for addressing this issue is to apply graph sparsification. The sparsification technique that we found to be effective for 2-HNC is the $k$-nearest neighbor (kNN) sparsification (Blum and Chawla, 2001) where samples $i$ and $j$ are connected only if $i$ is among the $k$ nearest neighbors of $j$, or vice versa. This results in a graph representation $G=(V,E)$ where $E$ is the set of pairs of samples where in each pair, one sample is one of the $k$ nearest neighbors of the other sample in the pair.
@@ 5 Implementation of 2-HNC
This section includes the specification of several implementation details. First, we give a brief description of the parametric cut solver used in this work. Second, we describe the choice of $k$ in the k-nearest neighbor graph sparsification method, as mentioned in the previous section. Finally, we quantify the pairwise similarity measure between pairs of samples. Our code implementation of 2-HNC can be found at https://github.com/hochbaumGroup/Positive-Unlabeled-Learning.
@@ 5.1 Parametric Cut Solver
[…]
@@ 6.4 Results on Synthetic Datasets
The histogram plots in Figures 4(a) and 4(c) show that, on average, 2-HNC gives positive improvement in terms of classification accuracy over uPU and PU ET, as the red dashed lines (the mean) lie to the right of the black solid line (the zero line). The average improvements, reported in Table 3, show the improvements of 2-HNC over uPU and PUET by $5.33\%$ and $8.02\%$, respectively. The fractions of datasets where 2-HNC improves on these two benchmarks are more than half, and in particular, as compared to PU ET, 2-HNC produces better results on almost all datasets ($92.96\%$). However, 2-HNC does not deliver consistent improvement in accuracy over nnPU, with the average improvement that is slightly below zero, at $0.44\%$, reported in Table 3 and shown in Figure 4(c) where the red dashed line (the mean) lies almost on the solid line (the zero line). The percentage of datasets where 2-HNC improves over nnPU is just above $50\%$.
@@ 6.5 Results on Real Datasets
This is also the case when we consider the balanced accuracy of 2-HNC and the best benchmark on MNIST where the p-value is $0.0711$, which is larger than $0.05$. Hence, it cannot be concluded with statistical significance that 2-HNC has a worse performance than nnPU. The underperformance of 2-HNC is statistically significant for its accuracy on Phishing and MNIST as compared to PU ET, in which the p-values are $0.0004$ and $3.3e\text{-}9$, respectively.